OpenStudy (anonymous):

Consider the equilibrium reaction: N2O4(g) yields 2NO2(g). Which of the following correctly describes the relationship between Kc and Kp for the reaction. A. Kp = Kc/RT B. Kp = (RT x Kc)-1 C. Kp = RT x Kc D. Kp = Kc E. Kp = RT/Kc

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

I'm not sure..)

OpenStudy (anonymous):

I thought it would be A because it relates to the original equation

OpenStudy (anonymous):

I thought because Kp = Kc (RT)^n n = 2 - 1 = 1

OpenStudy (anonymous):

So because of the coefficient subtracting from on another it change the overall equation to Kp=RT(Kc)

OpenStudy (anonymous):

i think so...you may've looked that eq. as Kc = Kp/(RT)^n both are the same..i mean Kc = Kp/(RT)^ and Kp = KcRT.. but again, I'm not sure) so after getting the right answer let me know.)

OpenStudy (anonymous):

ok I will. I guess I will keep the this question posted.